# Interplay between altermagnetism and topological superconductivity in an unconventional superconducting platform

### v1

@@ Abstract
We propose a theoretical model to investigate the interplay between altermagnetism and $p$-wave superconductivity, with a particular focus on topological phase transitions in a two-dimensional (2D) $p$-wave superconductor, considering both chiral and helical phases. Our study reveals that the emergence of helical and chiral Majorana states can be tuned by the amplitude of a $d-$wave altermagnetic order parameter, with the outcome depending on the nature of the superconducting state. In the helical superconductor, such an altermagnet can induce a topological phase transition into a gapless topological superconductor hosting Majorana flat edge modes (MFEMs). On the other hand, in the chiral superconductor, the topological transition takes place between a topologically nontrivial gapped phase and a gapless nodal-line superconductor, where the Bogoliubov quasiparticle bands intersect at an isolated line in momentum space. Remarkably, we show that when such an altermagnet is coupled to a mixed-pairing superconductor, with both chiral and helical components, a hybrid topological phase emerges, featuring dispersive Majorana edge that modes coexist with nearly flat Majorana edge states. Our findings therefore establish a novel platform for controlling and manipulating Majorana modes in unconventional superconductors with vanishing total magnetization.
@@ I Introduction
Altermagnetism, a recently established form of magnetic order, bridges the gap between ferromagnetism and antiferromagnetism [1, 2, 3]. While exhibiting zero net magnetization, a characteristic often associated with antiferromagnets, it simultaneously induces spin-split electronic band structures—a hallmark of ferromagnetism. This unique feature has sparked significant interest within both theoretical and experimental condensed matter communities [1, 2, 4, 3, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The intriguing nature of altermagnetism extends further since its textures manifest as higher-harmonic orders in momentum space (e.g., $d-$, flimit-$f-$, $g-$wave), thus providing the previously lacking magnetic analogs to unconventional superconducting (SC) orders, such as $p-$, $d-$, flimit-$f-$wave, etc. This opens exciting avenues for exploring novel material properties and functionalities, which requires further understanding of the underlying mechanisms governing altermagnetism and its potential applications.
@@
The realization of one- and two-dimensional (2D) topological superconductors has been a key impetus in the study of topological phases of matter, driven by the goal of creating localized Majorana zero modes, which are considered as fundamental building blocks for topological quantum computation [23, 24, 25, 26, 27, 28, 29, 30]. Furthermore, there has recently been growing interest in gapless topological superconductors from both theoretical and experimental perspectives [31, 32, 33, 34, 35, 36, 37, 38, 39], which may host zero-energy Majorana flat edge modes (MFEMs) instead of isolated Majorana zero modes, with a very recent experiment reporting the signature of MFEMs in a magnetic-superconductor hybrid system composed of Fe/Ta(110) [31]. We here present an alternative approach to introduce gapless topological superconductors hosting MFEMs, along with other new topological superconducting (TSC) phases, by invoking an unconventional SC platform involving an altermagnet, therefore carrying a net zero magnetization, that, to the best of our knowledge, has not been explored until now.
@@
To this end, we here consider the simplest $d$-wave altermagnetic order parameter (OP) in the presence of both chiral and helical $p$-wave SC states. We demonstrate that coupling the altermagnetic OP with this unconventional superconductor can give rise to several novel TSC phases. In the case of a helical $p$-wave superconductor, altermagnetism drives the system from a helical gapped topological superconductor to a gapless topological superconductor that hosts MFEMs. In contrast, for a chiral $p$-wave superconductor, a topological phase transition occurs, featuring the transition from a chiral gapped $p$-wave topological superconductor to a gapless nodal-line superconductor in the presence of altermagnetism. Notably, for a mixed $p$-wave superconductor, combining helical and chiral SC OPs, altermagnetism induces an unusual hybrid topological phase where both dispersive and flat Majorana edge modes coexist. Finally, this work reveals a tunable altermagnet-unconventional-superconductor platform that can support both gapless topological superconductivity, and a hybrid TSC phase that features coexisting dispersive and flat Majorana edge modes.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce the model for the 2D altermagnet with $d$-wave magnetic ordering coupled with unconventional helical and chiral $p$-wave superconductors. Next, we present numerical results on the edge states using the nanoribbon termination in the square lattice geometry in Sec. III, with the results for the edge state spectrum and the local density of states (LDOS) shown in Sec. III.1 and Sec. III.2, respectively, for the chiral and helical phases, while in Sec. III.3 we discuss both the edge states and LDOS in the case of mixed pairing phase. In Sec. IV, we compute the bulk gap profile as a signature of a topological phase transition by analyzing the corresponding lattice model. Finally, we summarize and discuss our results in Sec. V.
@@ II Model and Method
We here consider the Bogoliubov–de Gennes (BdG) Hamiltonian, representing an interplay between altermagnetism and 2D $p$-wave superconductor with helical and chiral order that can be written as,
@@
where the four-component Nambu spinor is $\Phi_{\bf k}=\left(c_{{\bf k}\uparrow},c_{{\bf k}\downarrow},c_{-{\bf k}\uparrow}^{\dagger},c_{-{\bf k}\downarrow}^{\dagger}\right)^{\top}$, where $c_{{\bf k}\uparrow}(c_{{\bf k}\downarrow}^{\dagger})$ is the annihilation (creation) operator of the quasiparticle at momentum ${\bf k}$ and spin up (momentum $-{\bf k}$ and spin down), with
$h_{\bf k}$ $=$ $\xi_{\bf k}\tau_{z}s_{0}+J({\bf k})\tau_{0}s_{x}+\Delta_{p}^{h}(\tau_{y}s_{0}\sin k_{x}-\tau_{x}s_{z}\sin k_{y})$ (2)
$+\Delta_{p}^{c}(\tau_{x}s_{x}\sin k_{x}-\tau_{y}s_{x}\sin k_{y}).$
Here, the electron quasiparticle dispersion is $\xi_{\bf k}=-2t\left(\cos k_{x}+\cos k_{y}\right)+4t-\mu$, with $\mu$ as the chemical potential. Furthermore, we assume that altermagnetic OP possesses $d$-wave symmetry,
$J({\bf k})=2J_{A}\left(\cos k_{x}-\cos k_{y}\right),$ (3)
with the amplitude of the altermagnetic order $J_{A}$ [14, 5, 2, 40]. The helical and chiral superconducting OPs are $\Delta_{p}^{h}$ and $\Delta_{p}^{c}$, respectively [41, 42, 43, 44, 45]. The Pauli matrices $\tau_{i}$ and $s_{i}$, $i=0,x,y,z$, act on particle-hole space and spin space, respectively. We obtain the spectrum corresponding to the helical superconductor, given by the Hamiltonian in Eq. (2), as
$\epsilon_{\bf k}=r\sqrt{\xi_{\bf k}^{2}+J({\bf k})^{2}-\frac{1}{2}(\Delta_{p}^{h})^{2}(\cos 2k_{x}+\cos 2k_{y})+sF({\bf k})}.$ (4)
For chiral superconductor, on the other hand, since the SC piece of the Hamiltonian commutes with the altermagnetic OP, the spectrum takes the form
@@
where $r,s=\pm$ are the band and particle-hole index, respectively. The form of $F({\bf k})$ reads as
$F({\bf k})=2J({\bf k})\sqrt{\xi_{\bf k}^{2}+(\Delta_{p}^{h}\sin k_{x})^{2}},$ (6)
and the topological phase transitions therefore take place for the critical strength of the altermagnetic OP,
$J^{c}_{A}=\pm\frac{1}{4}\left|4t-\mu\right|,$ (7)
with band gap closing at the high-symmetry points $X=(\pi,0)$ and vice versa.
We show the band structure [Eq. (4)] in Fig. 1 for a helical $p$-wave superconductor. Notably, a topological phase transition occurs at the amplitude of the altermagnetic OP $J_{A}=J_{A}^{c}$, as given by Eq. (7), separating two distinct topological states. For $J<J_{A}^{c}$, the system is always in a gapped, topologically non-trivial phase due to the presence of the helical $p$-wave superconductivity [see Fig.1(a)]. In contrast, for $J>J_{A}^{c}$, a new topologically non-trivial phase emerges, leading to gapless topological superconductivity [see Fig.1(c)], which hosts MFEMs. This new phase features an even number of gapless nodes [see Fig.1(d)], therefore representing a weak topological superconducting phase [46, 47], since it can be viewed as a stack of one-dimensional superconductors in class D (without time-reversal symmetry) changing the topological invariant at the (gapless) nodal points.
In Fig. 2, we display the band structure (using Eq. (5)) for a chiral $p$-wave superconductor. We observe a topological phase transition at $J_{A}=J_{A}^{c}$, where for $J_{A}<J_{A}^{c}$, the system is in a gapped, topologically non-trivial phase with broken time-reversal symmetry. On the other hand, for $J_{A}>J_{A}^{c}$, the system features a gapless nodal line superconductor (see Fig. 2(c)), where the closest to zero energy Bogoliubov quasiparticle bands intersect along a line in the Brillouin zone, with the corresponding form of the Fermi surface displayed in Fig. 2(d). Therefore, for a chiral $p$-wave superconductor, the $d$-wave symmetry of the altermagnet drives the system from a gapped topological phase to a gapless nodal line superconductor. In the following section, we discuss the signatures of these nodal superconductors in terms of the edge states.
@@ III Numerical Results
In this section, we discuss numerical calculations based on the lattice model implemented in real space given by the Hamiltonian in Eq. (1).
@@ III.1 Edge states in ribbon geometry
The numerical calculations of the edge states in the ribbon geometry are performed by employing periodic boundary conditions (PBC) along the $x$-direction and open boundary conditions (OBC) along the $y$-direction.
We first show the spectrum of the edge states for the helical $p$-wave superconductor for $J_{A}<J_{A}^{c}$ [$J_{A}>J_{A}^{c}$], in Fig. 3(a) [Fig. 3(b)]. For $J_{A}<J_{A}^{c}$, the system exhibits helical edge states protected by time-reversal symmetry, as shown in Fig. 3(a). On the other hand, the system leads to three flat edge modes for $J_{A}>J_{A}^{c}$ [shown in Fig. 3(b)], indicating six gapless points in the momentum space, as displayed in Fig. 1(d). This result closely parallels findings in the study of magnetic textures within hybrid superconductor systems [34], featuring the SC pairing with $d$-wave symmetry. Notably, in our context, the $d$-wave symmetry of the altermagnet plays a key role in facilitating gapless topological superconductivity, which is conducive to the emergence of MFEMs in helical $p$-wave superconductors.
In a chiral $p$-wave superconductor, the ribbon geometry, as expected, supports two chiral edge channels for a small magnitude of the altermagnet OP $J_{A}<J_{A}^{c}$, as displayed in Fig. 3(c). On the other hand, for the overcritical amplitude, $J_{A}>J_{A}^{c}$, the behavior changes dramatically as the edge modes become delocalized. Such a change in the band structure is a consequence of the transition to a nodal line superconducting phase, characterized by the presence of the isolated lines in the momentum space where the superconducting gap closes. The emergence of this phase is further corroborated by the plot of the corresponding (gapless) nodal lines in Fig. 2(d).
Figure 3: Edge states in the ribbon geometry for a helical and chiral $p$-wave superconductor coupled to the altermagnet [Eq. (2)]. (a) [(b)] Helical $p$-wave superconductor for the altermagnet coupling $J_{A}=0.3t$ [$J_{A}=0.65t$]. (c) [(d)] Chiral $p$-wave superconductor for the altermagnet coupling $J_{A}=0.3t$ [$J_{A}=0.8t$], respectively, for a chiral $p$-wave superconductor. We employ open [periodic] boundary conditions along $y$ [$x$] direction and $\mu=2t$ is fixed.

### v2

# Interplay between altermagnetism and topological superconductivity on an unconventional superconducting platform

@@ Abstract
We propose a theoretical model to investigate the interplay between altermagnetism and $p$-wave superconductivity, with a particular focus on topological phase transitions in a two-dimensional (2D) $p$-wave superconductor, considering both chiral and helical phases. Our study reveals that the emergence of helical and chiral Majorana states can be tuned by the amplitude of a $d$-wave altermagnetic order parameter, with the outcome depending on the nature of the superconducting state. In the helical superconductor, such an altermagnet can induce a topological phase transition into a gapless topological superconductor hosting Majorana flat edge modes. On the other hand, in the chiral superconductor, the topological transition takes place between a topologically nontrivial gapped phase and a gapless nodal-line superconductor, where the Bogoliubov quasiparticle bands intersect at an isolated line in momentum space. Remarkably, we show that when such an altermagnet is coupled to a mixed-pairing superconductor, with both chiral and helical components, a hybrid topological phase emerges, featuring dispersive Majorana edge modes that coexist with nearly flat Majorana edge states. Our findings therefore establish a novel platform for controlling and manipulating Majorana modes in unconventional superconductors with vanishing total magnetization.
@@ I Introduction
Altermagnetism, a recently established form of magnetic order, bridges the gap between ferromagnetism and antiferromagnetism [1, 2, 3]. While exhibiting zero net magnetization, a characteristic often associated with antiferromagnets, it simultaneously induces spin-split electronic band structures—a hallmark of ferromagnetism. This unique feature has sparked significant interest within both theoretical and experimental condensed matter communities, with works focusing on purely altermagnetic features [1, 2, 3, 4, 5, 6, 7], various superconductor responses to an altermagnetic state [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], and electronic topology in the presence of altermagnetism [24, 25]. The intriguing nature of altermagnetism extends further since its textures manifest as even-momentum higher-harmonic orders (e.g., $d$-, $g$-wave), thus providing the previously lacking magnetic analogs to unconventional superconducting (SC) orders, such as $d$-, $g$-wave, etc. This opens exciting avenues for exploring novel material properties and functionalities, which requires further understanding of the underlying mechanisms governing altermagnetism and its potential applications.
System PHS TRS Chiral Symmetry Symmetry Class Topological Invariant
Helical SC DIII $\mathbb{Z}_{2}$
Chiral SC D $\mathbb{Z}$
Helical SC + Altermagnet D $\mathbb{Z}$
Chiral SC + Altermagnet D $\mathbb{Z}$
Table 1: Topological classification of the Bogoliubov–de Gennes (BdG) Hamiltonian in Eq. (1) for different superconducting (SC) pairings, with and without coupling to the $d$-wave altermagnetic order parameter. Anti-unitary particle-hole symmetry (PHS) and time-reversal symmetry (TRS) operators square to +1+1 and 1-1, respectively, with the corresponding topological invariants determined according to the tenfold classification of topological insulators and superconductors [26]. The symbols ✓ and ✗ denote the presence and absence of the symmetry, respectively. See App. A for details of the symmetry transformations.
@@
The realization of one- and two-dimensional (2D) topological superconductors has been a key impetus in the study of topological phases of matter, driven by the goal of creating localized Majorana zero modes, which are considered as fundamental building blocks for topological quantum computation [27, 28, 29, 30, 31, 32, 33, 34]. Furthermore, there has recently been growing interest in gapless topological superconductors from both theoretical and experimental perspectives [35, 36, 37, 38, 39, 40, 41, 42, 43], which may host zero-energy Majorana flat edge modes (MFEMs) instead of isolated Majorana zero modes, with a very recent experiment reporting the signature of MFEMs in a magnetic-superconductor hybrid system composed of Fe/Ta(110) [35]. We here present an alternative approach to introduce gapless topological superconductors hosting MFEMs, along with other new topological SC phases, by invoking an unconventional SC platform involving an altermagnet, therefore carrying a net zero magnetization, that, to the best of our knowledge, has not been explored until now.
@@
To this end, we here consider the simplest $d$-wave altermagnetic order parameter (OP) in the presence of both chiral and helical $p$-wave SC states, as well as of the hybrid chiral-helical SC phase. We demonstrate that coupling such an altermagnetic OP featuring the spin oriented in the $x$-direction with this unconventional superconductor can give rise to several gapped and gapless topological SC phases. For a helical $p$-wave superconductor, altermagnetism drives the system from a helical gapped topological superconductor to a gapless topological superconductor that hosts MFEMs (Figs. 1 and 3). In contrast, for a chiral $p$-wave superconductor, a topological phase transition occurs from a chiral gapped $p$-wave topological superconductor to a gapless nodal-line superconductor in the presence of altermagnetism (Figs. 2 and 4). Notably, for a mixed $p$-wave superconductor, combining helical and chiral SC OPs, altermagnetism induces an unusual hybrid topological phase where both dispersive and flat Majorana edge modes coexist, as displayed in Fig. 5. In all the discussed cases, the topological phases are separated by the bulk band closing, as explicitly shown in Fig. 6. Finally, this work reveals a tunable altermagnet-unconventional-superconductor platform that can support both gapless topological superconductivity, and a hybrid topological SC phase that features coexisting dispersive and flat Majorana edge modes.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce the model for the 2D altermagnet with $d$-wave magnetic ordering coupled with unconventional helical and chiral $p$-wave superconductors. Next, we present numerical results on the edge states using the nanoribbon termination in the square lattice geometry in Sec. III, with the results for the edge state spectrum and the local density of states (LDOS) shown in Sec. III.1 and Sec. III.2, respectively, for the chiral and helical phases, while in Sec. III.3 we discuss both the edge states and LDOS in the case of mixed pairing phase. In Sec. IV, we compute the bulk gap profile as a signature of a topological phase transition by analyzing the corresponding lattice model. Finally, we summarize and discuss our results in Sec. V. Analysis of symmetries of the SC Hamiltonian is presented in App. A, while the details of the edge theory in purely helical SC phase are discussed in App. B.
@@ II Model and Method
In this work, we analyze a minimal model that captures the nontrivial interplay between altermagnetism and 2D unconventional superconductivity. Specifically, we consider a Bogoliubov–de Gennes (BdG) Hamiltonian that incorporates both helical and chiral $p$-wave SC orders in the presence of an altermagnetic background. The full BdG Hamiltonian describing this hybrid system reads
@@
where $\Phi_{\bf k}=\left(c_{{\bf k}\uparrow},c_{{\bf k}\downarrow},c_{-{\bf k}\uparrow}^{\dagger},c_{-{\bf k}\downarrow}^{\dagger}\right)^{\top}$ is the four-component Nambu spinor. Here, $c_{{\bf k}\uparrow}$ annihilates an electron quasiparticle with momentum $\bf k$} and spin $\uparrow$, while $c_{-{\bf k}\downarrow}^{\dagger}$ creates a quasiparticle with opposite momentum and spin. The BdG kernel $h_{\bf k}$ takes the form
$h_{\bf k}$ $=$ $\xi_{\bf k}\tau_{z}s_{0}+J({\bf k})\tau_{0}s_{x}+\Delta_{p}^{h}\left(\tau_{y}s_{0}\sin k_{x}-\tau_{x}s_{z}\sin k_{y}\right)$ (2)
$+\Delta_{p}^{c}\left(\tau_{x}s_{0}\sin k_{x}-\tau_{y}s_{0}\sin k_{y}\right),$
where $\tau_{i}$ and $s_{i}$ ($i=0,x,y,z$) are Pauli matrices acting in particle-hole (Nambu) and spin space, respectively, and the free-quasiparticle dispersion is given by
$\xi_{\bf k}=-2t(\cos k_{x}+\cos k_{y})+4t-\mu,$ (3)
where $t$ is the nearest-neighbor hopping amplitude and $\mu$ is the chemical potential.
The term $\sim J({\bf k})$ encodes the coupling of the altermagnetic order to the electron quasiparticles, which we assume to possess $d$-wave symmetry,
$J({\bf k})=2J_{A}(\cos k_{x}-\cos k_{y}),$ (4)
where $J_{A}$ is the amplitude of the altermagnetic exchange field [13, 10, 2, 44]. The momentum-dependent structure of the altermagnetic OP leads to spin-split quasiparticle bands that break fourfold rotational symmetry ($C_{4}$), $J({\bf k})\rightarrow-J({\bf k})$ under $C_{4}$. In fact, this antisymmetry under lattice rotations distinguishes altermagnetism from conventional ferromagnetic exchange, which is momentum-independent. Moreover, despite inducing spin splitting, the altermagnetic order preserves zero net magnetization over the Brillouin zone, as the contributions from different momenta cancel out. The altemagnetic term also breaks time-reversal symmetry (TRS) explicitly, since under the time-reversal operation $\mathcal{T}$, one has $s_{x}\rightarrow-s_{x}$ while $J({\bf k})$ remains invariant, such that $\mathcal{T}[J({\bf k})s_{x}]\mathcal{T}^{-1}=-J({\bf k})s_{x}$. This results in a nontrivial coupling to the SC sector, modifying the quasiparticle spectrum in a momentum-anisotropic and symmetry-dependent fashion, which, in turn, also pertains to the topological features of the Bogoliubov band structures in the resulting SC state, as we show in the following.
The two SC OPs considered in this model correspond to different pairing symmetries. The helical $p$-wave OP, $\Delta_{p}^{h}$, features a spin-triplet pairing with a spin-dependent orbital structure preserving TRS, while the chiral $p$-wave OP, $\Delta_{p}^{c}$, breaks time-reversal, see also Table 1. A more detailed discussion on the symmetries is deferred to App. A.
We now analyze the quasiparticle dispersion and the band structure of the superconductor-altermagnet system for both pairing symmetries. In the helical case, the energy spectrum takes the form
$\epsilon_{\bf k}=r\sqrt{\xi_{\bf k}^{2}+J({\bf k})^{2}-\frac{1}{2}(\Delta_{p}^{h})^{2}(\cos 2k_{x}-\cos 2k_{y})+sF({\bf k})},$ (5)
where $r,s=\pm$ label the band and particle-hole indices, respectively, and
$F({\bf k})=2J({\bf k})\sqrt{\xi_{\bf k}^{2}+(\Delta_{p}^{h}\sin k_{x})^{2}}.$ (6)
The term $F({\bf k})$ captures the momentum-resolved coupling between the helical pairing and the altermagnetic field, leading to anisotropic features in the spectrum that can drive topological transitions.
In contrast, for the chiral pairing, the SC term commutes with the altermagnetic OP, and the spectrum simplifies to
@@
with the altermagnetic piece therefore shifting the spectrum of the Bogoliubov quasiparticles albeit in a momentum-dependent manner, which, in turn, can also yield topologically nontrivial features in the resulting bands.
Importantly, the interplay between altermagnetism and superconductivity enables topological phase transitions, signaled by the closing and reopening of the quasiparticle gap at high-symmetry points in the Brillouin zone. For the helical case, the critical value of the altermagnetic coupling $J_{A}$ at which such a transition occurs is found by locating the band-touching point at $X=(\pi,0)$,
$J^{c}_{A}=\pm\frac{1}{4}|4t-\mu|,$ (8)
which identifies a topological phase boundary in parameter space.
We show the band structure [Eq. (5)] in Fig. 1 for a helical $p$-wave superconductor. Notably, a topological phase transition occurs at the amplitude of the altermagnetic OP $J_{A}=J_{A}^{c}$, as given by Eq. (8), separating two distinct topological states. For $J<J_{A}^{c}$, the system is always in a gapped, topologically non-trivial phase due to the presence of the helical $p$-wave superconductivity [see Fig.1(a)]. In contrast, for $J>J_{A}^{c}$, a new topologically non-trivial phase emerges, leading to gapless topological superconductivity [see Fig.1(c)], which hosts MFEMs. This new phase features an even number of gapless nodes [see Fig.1(d)], therefore representing a weak topological SC phase [45, 46], since it can be viewed as a stack of one-dimensional superconductors in class D (without TRS), see also Table 1, changing the topological invariant at the (gapless) nodal points.
Figure 3: Edge states in the ribbon geometry for a helical and chiral $p$-wave superconductor coupled to the altermagnet [Eq. (2)]. (a) [(b)] Helical $p$-wave superconductor for the altermagnet coupling $J_{A}=0.3t$ $J_{A}=0.65t$]. (c) [(d)] Chiral $p$-wave superconductor for the altermagnet coupling $J_{A}=0.3t$ $J_{A}=0.8t$], respectively, for a chiral $p$-wave superconductor. We employ open [periodic] boundary conditions along $y$ [$x$] direction and $\mu=2t$ is fixed.
Figure 4: Panels (a) and (b) show local density of states (LDOS) at zero energy ($E=0$) of the helical $p$-wave superconductor coupled to the altermagnet (Eq. (1)) for the values of the altermagnet amplitude $J_{A}=0.3t$ and $J_{A}=0.65t$, respectively. Panels (c) and (d) display LDOS at $E=0$ of the chiral $p$-wave superconductor for the coupling constant $J_{A}=0.3t$ and $J_{A}=0.8t$, respectively. The inset in each panel shows the spectrum of the closest-to-zero-energy states in the corresponding phase, with the $E=0$ states being of Majorana nature, as corroborated by the characteristic equal distribution of the spectral weight between the particle and hole components of the corresponding wavefunctions. We here set $\mu=2t$ and $\Delta_{p}^{h}=\Delta_{p}^{c}=1.0$, and $t=1$.
In Fig. 2, we display the band structure [using Eq. (7)] for a chiral $p$-wave superconductor. We observe a topological phase transition at $J_{A}=J_{A}^{c}$, where for $J_{A}<J_{A}^{c}$, the system is in a gapped, topologically non-trivial phase with broken TRS. On the other hand, for $J_{A}>J_{A}^{c}$, the system features a gapless nodal line superconductor [see Fig. 2(c)], where the closest to zero energy Bogoliubov quasiparticle bands intersect along a line in the Brillouin zone, with the corresponding form of the Fermi surface displayed in Fig. 2(d). Therefore, for a chiral $p$-wave superconductor, the $d$-wave symmetry of the altermagnet drives the system from a gapped topological phase to a gapless nodal line superconductor. In the following section, we discuss the signatures of these nodal superconductors in terms of the edge states.
@@ III Numerical Results
In this section, we discuss numerical calculations based on the lattice model implemented in real space given by the Hamiltonian in Eq. (1). The numerical analysis of edge states and local density of states (LDOS) presented in this section reveals that the interplay between d-wave altermagnetism and unconventional superconductivity gives rise to a range of nontrivial topological phenomena, particularly, it captures phase transitions and the emergence of Majorana modes across different SC regimes.
@@ III.1 Edge states in ribbon geometry
The edge state spectrum in the ribbon geometry is obtained through numerical computations that are carried out by imposing periodic boundary conditions (PBC) along the $y$-direction, implying the conserved $k_{y}$ momentum along the periodic edge, and open boundary conditions (OBC) along the $x$-direction.
We first show the spectrum of the edge states for the helical $p$-wave superconductor in the two regimes: $J_{A}<J_{A}^{c}$ [see Fig.3(a)] and $J_{A}>J_{A}^{c}$ [see Fig.3(b)]. For $J_{A}<J_{A}^{c}$, the system exhibits robust helical edge states, which, however, cease to be protected by TRS due to the presence of the coupling to the altermagnet, see Table 1 and Appendix A. The loss of TRS protection renders the helical edge states vulnerable to backscattering. Nevertheless, these edge modes can be directly traced back to the helical modes that exist in the absence of any coupling to the altermagnet. Thus, their presence reflects the underlying topological character of the parent helical $p$-wave SC phase, albeit with modified symmetry protection due to the finite $J_{A}$.
On the other hand, the system leads to three flat edge modes for $J_{A}>J_{A}^{c}$ [shown in Fig. 3(b)], indicating six gapless points in the momentum space, as displayed in Fig. 1(d), and therefore the system enters a nodal superconductor phase. This result closely parallels findings in the study of magnetic textures within hybrid superconductor systems [38], featuring the SC pairing with $d$-wave symmetry. Notably, in our context, the $d$-wave symmetry of the altermagnet plays a key role in promoting gapless topological superconductivity, which leads to the emergence of MFEMs in helical $p$-wave superconductors.
In a chiral $p$-wave superconductor, the ribbon geometry, as expected, supports two chiral edge channels for a small magnitude of the altermagnet OP, $J_{A}<J_{A}^{c}$, as displayed in Fig. 3(c), with the crossing points moved from the zero momentum since the altermagnetic cooupling commutes with the chiral pairing term. These edge channels are indicative of a topological SC phase gapped in the bulk that preserves a nontrivial Chern number, and they are localized at opposite edges of the system, reflecting the underlying broken TRS of the chiral pairing. On the other hand, for an overcritical amplitude of the altermagnetic coupling, $J_{A}>J_{A}^{c}$, the behavior of the system changes dramatically. The edge modes that were previously sharply confined to the boundaries become delocalized, which is a direct consequence of the system undergoing a topological phase transition into a gapless nodal-line SC phase, wherein the bulk gap closes along isolated lines in momentum space. The nodal-line phase is further confirmed by the gapless contours in the Brillouin zone shown in Fig. 2(d). Finally, this transition is also consequential for the local density of states, which will be discussed next.